Bermuda Cruises issues only common stock and coupon bonds. The firm has a debt–equity ratio of 1.25. The cost of equity is 12.7 percent and the pretax cost of debt is 7.2 percent. What is the capital structure weight of the firm's equity if the firm's tax rate is 39 percent?

Multiple Choice

  • .5162

  • .5556

  • .4444

  • .4166

  • .4897
Answer #1

Weight of equity = 1 / (1 + D-E)

Weight of equity = 1 / (1 + 1.25)

Weight of equity = 1 / 2.25

Weight of equity = 0.4444